OpenStudy (megannicole51):

Find the first three nonzero terms of the taylor series for e^x + e^(-x)... I know the expansion of both terms but im not sure how to add them to find the expansion!

OpenStudy (megannicole51):

@ganeshie8

ganeshie8 (ganeshie8):

put the both expansions first

ganeshie8 (ganeshie8):

btw, you expanding it about 0 right ?

OpenStudy (megannicole51):

e^x=1+x+x^2/2!+x^3/3!+.... e^(-x)=1-x+x^2/2!-x^3/3!+....

ganeshie8 (ganeshie8):

yes, clearly all odd powers cancel out.. . so simply add the remaining even powers

ganeshie8 (ganeshie8):

e^x = 1+x+x^2/2!+x^3/3!+.... e^(-x) = 1-x+x^2/2!-x^3/3!+.... -------------------------------------------------------------------- e^x + e^(-x) = 2 + 2*x^2/2! + 2*x^4/4! + .....

OpenStudy (megannicole51):

oh okay that makes sense!:) thank you!

ganeshie8 (ganeshie8):

np :)